I do not know how to solve this problem:
Convert 0.90 kg/cm^3 to mg/mm^3
PLEASE HELP.

Answer by htmentor(1343)

Convert 0.90 kg/cm^3 to mg/mm^3
======================================
There are 1000 mg in 1 g, 1000 g in 1 kg, and 10 mm in 1 cm, so we can write:
0.9 kg/cm^3 * 1000 mg/g * 1000 g/kg * 1 cm^3/(10 mm)^3 = 0.9*1000^2/10^3 = 900 mg/mm^3